find the equation of the line with a y-intercept of (0,7) that is perpendicular to y= 1/2x -4
Answer by stanbon(75887)

find the equation of the line with a y-intercept of (0,7) that is perpendicular to y= 1/2x -4
-----
The given equation has slope = 1/2
-----
Any line perpendicular to it has slope = -2
-------
Form: y = mx + b
---
You have slope = -2 and y = 7 when x = 0
7 = -2*0 + b
b = 7
-------
Equation:
y = -2x +7
==========================
Cheers,
Stan H.
======================